A manufacturer sells watches for $50 per unit. The fixed cost related to this product are $10,000 per month, and the variable cost per unit is $30.
a) Write the equations of the revenue and cost functions.
b) How many watches must be sold to break even?

Answer by addingup(3677)

10,000+30x = total cost (fixed+variable)
50x-(10,000+30x) = revenue
Breakeven:
10,000+30x = 50x
10,000 = 20x flip the equation to get the variable on the left (looks better):
20x = 10,000
x = 10,000/20 = 500